OpenStudy (anonymous):

Simplify by dividing three over five divided by negative four over nine.. negative twenty-seven over twenty. twenty-seven over twenty. negative twelve over forty-five. twelve over forty-five.

OpenStudy (jdoe0001):

\(\bf \cfrac{3}{5}\div\cfrac{4}{9}\implies \cfrac{3}{5}\cdot \cfrac{9}{4}\implies \cfrac{3\cdot 9}{5\cdot 4}\)

OpenStudy (sparklestaraa):

what he said is correct

OpenStudy (anonymous):

Ok so its b

OpenStudy (anonymous):

@Sparklestaraa

OpenStudy (sparklestaraa):

Yes i would think :)

OpenStudy (anonymous):

could you help with 1 more?

OpenStudy (sparklestaraa):

Yes sure

OpenStudy (anonymous):

Pedro simplified four over seven divided by five over eight.; his work is shown below. Identify where he made his error. Original division problem: four over seven divided by five over eight. Step 1: four over seven multiplied by eight over five. Step 2: four times five over seven times eight. Step 3: twenty over fifty-six. Step 4: five over fourteen. Step 1 Step 2 Step 3 Step 4

OpenStudy (anonymous):

@mathacomplice @mathhelper @parissfrance09

OpenStudy (sparklestaraa):

@campbell_st @Nnesha @pooja195 Help i think its Step 2

pooja195 (pooja195):

they need to repost this....